OpenStudy (anonymous):

Write the slope-intercept form of the equation for the line passing through the given pair of points. (-9,7) and (-7,4)

OpenStudy (anonymous):

y = mx + b

OpenStudy (vijay):

2y=-3x + 85

OpenStudy (anonymous):

First find the slope, then use point-slope form to write an equation. Slope: \[m = \frac{y_2 - y_1}{x_2 - x_1}\]Point-Slope Form:\[y - y_1 = m(x - x_1)\]